A buffered solution of 25 mL of an acid-base indicator is prepared by mixing 5mL of a stock solution of indicator 2,10x10^-4 M + 2,5 mL of Na2CO3 0,1 M + 10mL of NaHCO3 0,2 M
The stoichiometric equilibrium constant of carbonic acid are 4,47x10^-7 and 4,68x10^-11
Calculate the pH of this solution.
1
Expert's answer
2020-06-25T07:43:57-0400

The acid-base equilibrium: HCO3- = H+ + CO32-

Constant is 4.68*10-11 = [H+]*[CO32-]/[HCO3-]

Concentrations in resulting solution are:

[CO32-] = 0.1*2.5/(5+2.5+10) = ‭0.014 M

[HCO3-] = 0.2*10/(+2.5+10) = ‭0.114 M

Therefore: [H+] = 4.68*10-11 * 0.114 / 0.014 = 3.8*10-10 M

pH = -log[H+] = 9.42